OpenStudy (anonymous):

7 / x - 4 = -4 x / x +1

OpenStudy (anonymous):

@precal

OpenStudy (anonymous):

@wio @Michele_Laino

OpenStudy (michele_laino):

is your equation this? \[\frac{ 7 }{ x }-4=-4\frac{ x }{ x+1 }\]

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

ok! Now do you know what is the least common multiple?

OpenStudy (anonymous):

no

OpenStudy (michele_laino):

your least common multiple is: \[x*(x+1)\]

OpenStudy (michele_laino):

so yopu have to multiply both sides of your equation by x(x+1)

OpenStudy (michele_laino):

for example, left side is: \[x(x+1)*\left( \frac{ 7 }{ x }-4 \right)=x(x+1)*\frac{ 7 }{ x }+x(x+1)*(-4)=...\] please continue

OpenStudy (michele_laino):